# Liquid Resistance Liquid Capacitance Networks

### v2

@@ 4.1 Accuracy and Speed of LRCUs on RNNs Benchmarks
To achieve a somewhat similar number of parameters, we use 64 cells for LRCUs, and 100 cells for the others. The total number of parameters in a model also depends on the number of inputs and outputs of the tasks considered. Finally, we also conduct experiments in a relatively complex and high-dimensional, Image-based Regression Task, for lane keeping in autonomous vehicles.
All experimental details, including the hyperparameters and the number of trainable parameters, are given in Appendix A.3. In Table 1 we provide the accuracy of the various models compared. Moreover, in Figure 5 we give the validation accuracy convergence curves for the compared models.
Table 1: Classification accuracy in percentage. For similar number of trainable parameters, we used 64 cells in LRCUs and 100 cells in the other models. We repeated the experiments 5×5$\times$5 × for person activity, and 3×3$\times$3 × for IMDB and permuted sequential MNIST classification tasks, respectively.
Task LSTM GRU MGU LRCU-A LRCU-S
Localization $82.90\pm 0.31$ $82.76\pm 0.41$ $83.35\pm 0.30$ $83.90\pm 0.34$ $\mathbf{84.21\pm 0.26}$
IMDB $86.56\pm 0.49$ $86.32\pm 0.51$ $85.18\pm 0.85$ $\mathbf{87.00\pm 0.53}$ $85.73\pm 0.41$
psMNIST $91.20\pm 0.10$ $90.20\pm 0.44$ $87.78\pm 0.82$ $91.31\pm 0.33$ $\mathbf{91.74\pm 0.41}$
Figure 5: Accuracy convergence. Left: Irregularly sampled person activity dataset. As one can see, LRCUs converge a bit slower, yet they achieve the best accuracy after 100 epochs. Middle: IMDB sentiment classification problem. LRCU-A achieves the highest accuracy after only 5 epochs. The accuracy though, decreases a bit over time. Right: Permuted sequential MNIST dataset. LRCUs demonstrate a sligtly better validation accuracy when compared to GRUs and LSTMs.
@@
The localization data for the Person Activity dataset given in Vidulin et al. (2010), captures the recordings of five individuals, engaging in various activities. Each wore four sensors at the left and the right ankle, at the chest, and at the belt, while repeating the same activity five times. The associated task, is to classify their activity based on the irregularly sampled time-series. This task is a classification problem, adapted from Lechner and Hasani (2020).
In Table 1, we present the experimental results for the accuracy of the classification for LSTMs, GRUs, MGUs, and LRCUs. The experiments for the LRCUs were done with both the asymmetric (LRCU-A) and the symmetric (LRCU-S) elastance, respectively. As one can see, both LRCUs considerably outperformed the traditional gated RNNs. In particular, the LRCU-S performed best, achieving an accuracy of 84.99%, slightly better than the LRCU-A. In Figure 5 we compare the validation accuracy. LRCUs converge slower, but they both surpass the other models after 100 epochs.
@@
The IMDB Movie-Review Sentiment Classification dataset, also known as the Large Movie Review Dataset (Maas et al., 2011), is designed for binary sentiment classification. This data set includes 25,000 movie reviews, for both training and testing. Each review was labeled with either a positive or a negative sentiment.
Table 1 presents our experimental results for this task, too. While the LRCU-S achieved a performance which was comparable to the one of the traditional RNNs, the LRCU-A had the best performance, by achieving an accuracy of 87%. Moreover, LRCU-A achieves the highest accuracy after only 5 epochs, even though this decreases a bit over time, as can be seen in the second image of Figure 5.
@@
The Permuted Sequential MNIST dataset is a variant of the MNIST Digits Classification data set, designed to evaluate recurrent neural networks, adapted from Le et al. (2015). In this task, the 784 pixels (originally images of size $28{\times}28$) of digits are presented sequentially to the network. The main challenge lies in predicting the digit category, only after all pixels are observed. This task tests the network’s ability to handle long-range dependencies. To make the task even more complex, a fixed random permutation of the pixels is first applied.
Our experimental results are shown in Table 1. As before, both LRCUs surpass the accuracy of the other models. In particular, the LRCU-S achieves the best results. In Figure 5, we show that LRCUs achieve a slightly better validation accuracy when compared to GRUs and LSTMs on this task.
Figure 6: Lane-Keeping task. The first two figures show summer and winter conditions, respectively. The red rectangle indicates the input and the blue line the predicted steering angle of the network. The third figure illustrates the overall network architecture for this task. The CNN head extracts the input features of a video stream. They are passed to the recurrent policy, responsible for steering.
@@
LTC $0.204\pm 0.004$ $\mathbf{0.011\pm 0.0004}$
STC $0.201\pm 0.006$ $\mathbf{0.011\pm 0.0010}$
LRCU-A $0.154\pm 0.005$ $\mathbf{0.010\pm 0.0003}$
LRCU-S $0.142\pm 0.004$ $\mathbf{0.011\pm 0.0010}$
Figure 7: Accuracy results for the Lane-Keeping task, averaged over 3 seeds.
Lane-Keeping Task.
The dataset for the Lane-Keeping Task task is obtained from human-driving recordings captured under various weather and illumination conditions (Lechner et al., 2022), as shown in Figure 6. The network architecture contains a CNN-head for extracting the main features from the camera input. They are fed into the gated recurrent models, for the sequential-regression prediction. This network architecture is illustrated on the right in Figure 6. This setup is adapted from Farsang et al. (2023).
In Figure 7, we report the experimental losses of the models considered on the Lane-Keeping task. As it was the case before, both LRCU-A and LRCU-S obtain comparable validation losses, with LRCU-A achieving the best weighted-validation loss (loss weighted by road curvature). This is arguably the more important loss. The validation loss curve is presented on the right side in Figure 7.
4.2 Accuracy and Speed of LRCs on Neural ODEs Benchmarks
LRCs are competitive in accuracy and convergence speed on popular Neural-ODE benchmarks. As two examples, we consider a linear Spiral towards the origin and the non-linear Duffing oscillator.
In our experiments we use an LRC with 16 hidden states connected all to all. LRCs are highly suitable for such tasks as they are ODEs. We also use two extra layers for mapping. The former maps the state of these systems to the dimensionality of the hidden state, allowing us to learn more expressive 16-dimensional dynamics. The latter remaps these dynamics back to the original system state dimensionality. As one can see in Figure 8, we obtain for both tasks an arguably perfect fit.
We compare the performance of the LRCs with a 3-layer Neural-ODE, containing 32, 32, and 2 neurons, respectively. These three layers match the number of trainable parameters, with the ones of the LRCs. Neural-ODEs are solved by Runge-Kutta 4(5) of Dormand-Prince, and LRCs by explicit Euler of order 1. The results of our experiments, averaged over 3 seeds, are shown in Figure 9. In the linear ODE task, LRCs converge slower but achieve a smaller loss compared to the Neural-ODE. In the nonlinear ODE task, LRCs are clearly able to match the oscillatory dynamics more effectively.
Figure 8: Neural-ODE experiments. Linear spiral (first two figures, from left to right) and non-linear oscillatory systems (last two figures, from left to right) with $(x,y)$ state variables, solved by LRCs.
@@ 4.2 Accuracy and Speed of LRCs on Neural ODEs Benchmarks
Linear $0.012\pm 0.005$ $\mathbf{0.009\pm 0.004}$
Non-linear $0.142\pm 0.117$ $\mathbf{0.003\pm 0.001}$
Figure 9: Loss convergence for the linear and non-linear ODEs is depicted in the first and second pictures, respectively. In both cases, LRCs converged to a smaller loss, indicating their ability to fit these dynamics better than the vanilla Neural-ODE. Test loss is shown in the table on the right.
@@ 4.3 Interpretability of LRCUs on the Autonomous-Driving Lane-Keeping Task
To show that LRCUs outperform LSTMs, GRUs, and MGUs in terms of interpretability, we conduct Lane-Keeping Task experiments, and compute the following interpretability metrics: the network attention, the absolute correlation of neural activity with road trajectory, and the activity of neurons while driving. In the Appendix we compute attention robustness by the structural similarity index.
Understanding neural-network attention during decision-making is a crucial element in increasing trust in the network. By using the VisualBackprop (Bojarski et al., 2016) method, we visualize which pixels in the input image have the most impact at the given timestep. Even though the same structure of CNN heads is used in the training pipeline, the different recurrent parts of the decision-making policy, have a big influence on the learned features in the convolutional part through backpropagation.
Figure 10 shows our results in the summer season, where lighter-highlighted regions indicate the attention of the network. We found that the LSTM takes into account irrelevant parts of the image, during its decision-making. The rest of the models have most of the attention on the road, and LRCU networks especially, focus on the horizon. The attention in winter season is given in the Appendix.
As another interpretability metric, we assess how the neural activity within a Lane-Keeping RNN-policy changes during deployment in the closed-loop simulation. Specifically, we are interested in identifying neurons exhibiting an activity matching the geometry of the trajectory. We conduct tests for a 1 km long drive, in both summer and winter. In Figure 11, we illustrate the activity of States $h_{1}$-$h_{2}$, corresponding to Cells 1-2 in the RNN policy, for summer, respectively. In the Appendix, we illustrate the activity of Cells 1-2 for winter. Our results reveal that the LSTM, GRU, MGU, and LRCU-A cells, have a scattered, hard to interpret activity, respectively. In contrast, LRCU-S cells demonstrate a very smooth, continuously changing activity, that aligns very well with the road’s trajectory. This is most likely due to their double-liquid resistance and capacitance time constant.
Table 2: Interpretability. Absolute correlation between the neural activity and the trajectory of the road. Values closer to 1 indicate stronger correlation. Results are averaged over 3 runs.
LSTM GRU MGU LRCU-A LRCU-S
Summer $0.400\pm 0.299$ $0.363\pm 0.256$ $0.317\pm 0.227$ $0.500\pm 0.314$ $\mathbf{0.645\pm 0.321}$
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Interpretability. The neural activity of two command cells in the learned policy for the Lane-Keeping Task, projected over time on the 1km road driven in summer, for three popular gated recurrent units (LSTMs, GRUs, and MGUs) and for LRC units (LRCU-A, and LRCU-S). It is very hard to visually match the neural activity of LSTM, GRU, and MGU cells to the traversed road geometry, respectively. However, LRCU-S cells especially, display a very smooth and identifiable pattern during driving: Cell 1 is responsible for turning right, and Cell 2 for turning left.
@@ 5 Discussion, Scope and Conclusion
We introduced Liquid-Resistance Liquid-Conductance networks (LRCs), a Neural ODE model which is more accurate, more efficient, and biologically more plausible, compared to neuroscience’s Electric Equivalent Circuits (EECs), and (saturated) Liquid Time-Constant networks (STCs and LTCs).
Due to their smooth behavior LRCs are easy to integrate and applicable to a wide range of scenarios, including continuous processes and datasets sampled regularly or irregularly. LRCs thus represent a promising advancement in the field of bio-inspired deep learning. LRCs integrated with explicit Euler of order 1, and a time-step of 1 (LRCUs), have a very similar structure to gated RNNs. They not only demonstrate better performance compared to LSTMs, GRUs, and MGUs on several benchmarks, but also exhibit significantly higher interpretability, demonstrated in the Lane-Keeping task.
In our experiments, we compared models with identical hyperparameters to ensure a fair evaluation, and isolating differences in the models themselves under standardized conditions. However, further hyperparameter tuning could lead to better performance. This limitation underscores the potential for optimization beyond the scope of our current study, suggesting opportunities for future research to explore and enhance model performance. Another limitation of our study is the scope of the benchmarks used, which could be larger in the future. We focused on analyzing our advanced bio-inspired model’s properties and showcasing its performance covering various scenarios.
Acknowledgements
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101034277.
Appendix A Appendix
This appendix provides the proofs of the two theorems stated for LCRs, discusses the RNN form of LRCUs and their connection to GRUs and other gated recurrent units, and gives additional details for the experimental evaluation, demonstrating the accuracy, efficiency, and interpretability of LRCs.
A.1 Proofs of the Smoothness and Accuracy Theorems
In this section we provide a detailed proof of the Smoothness and Accuracy Theorems stated in Section 3. The Smoothness Theorem first derives the local Lipschitz constant of LRCs and then shows that it is smaller than the one of STCs. The Accuracy Theorem, first computes the error of LRCs on the validation time series, and then shows that it is less than the one of the STCs.
A.1.1 Proof of the Smoothness Theorem
Theorem 1 (LRC Smoothness).
@@ A.1.1 Proof of the Smoothness Theorem
We now want to prove that there exists a choice of parameters for $\epsilon(w)$, such that the Lipschitz constant of the output $o^{r}$ of the LCR with respect to its hidden states and input $y=[h,x]$, is less than the one of $o^{s}$ in the STC. So the goal is to state conditions on the parameters of $\epsilon(w)$ such that:
$\max_{y}\|\nabla_{y}o^{s}\|_{2}\leq\lambda^{s}\ \wedge\max_{y}\|\nabla_{y}o^{r}\|_{2}\leq\lambda^{r}\quad\Rightarrow\quad\lambda^{r}<\lambda^{s}$ (6)
@@
$o=Qh,\quad$ $\quad o_{k}=\sum_{i=1}^{m}q_{ki}h_{i}$ (7)
$\partialderivative{o_{k}}{y_{j}}$ $=\sum_{i=1}^{m}q_{ki}\partialderivative{h_{i}}{y_{j}}$ (8)
@@
[…]
$\leq 1+\left|\int_{t-\Delta t}^{t}\max_{y}\partialderivative{\dot{h}}{y_{t,j}}dt\right|$ (11)
@@
$\sigma^{\prime}(x)\leq 0.25,$ $\quad\tau^{\prime}(x)\leq 1$ (15)
Using these bounds, for the symmetric version of elastance in Equation (5) it holds that:
$\epsilon^{\prime}(w_{i})=\sigma^{\prime}(w_{i}+k_{i})-\sigma^{\prime}(w_{i}-k_{i})\leq\max\sigma^{\prime}(w_{i}+k_{i})-\min\sigma^{\prime}(w_{i}-k_{i})\leq 0.25$ (16)
@@
$\frac{\epsilon^{\prime}(w_{i})}{1-\epsilon(w_{i})}\leq\epsilon(w_{i})$ (17)
For asymmetric $\epsilon$ this is straightforward as the derivivative of a sigmoid is $\sigma^{\prime}(x)=\sigma(x)\cdot(1-\sigma(x))$. For symmetric $\epsilon$ we need to reformulate the equations using that $\sigma(w_{i}+k_{i})\geq\sigma(w_{i}-k_{i})$ for $k_{i}\geq 0$:
$\frac{\epsilon^{\prime}(w_{i})}{1-\epsilon(w_{i})}$ $=\frac{\sigma(w_{i}+k_{i})(1-\sigma(w_{i}+k_{i}))-\sigma(w_{i}-k_{i})(1-\sigma(w_{i}-k_{i}))}{1-\sigma(w_{i}+k_{i})+\sigma(w_{i}-k_{i})}$ (18)
$\leq\frac{\sigma(w_{i}+k_{i})(1-\sigma(w_{i}+k_{i}))-\sigma(w_{i}-k_{i})(1-\sigma(w_{i}+k_{i}))}{1-\sigma(w_{i}+k_{i})}$ (19)
$=\epsilon(w_{i})$ (20)
@@
$\leq|\sigma^{\prime}(f_{i})g_{ji}\sigma^{\prime}(a_{ji}y_{j}+b_{ji})a_{ji}h^{s}_{i}|+|\tau^{\prime}(u_{i})k_{ji}\sigma^{\prime}(a_{ji}y_{j}+b_{ji})a_{ji}e_{li}|+1$ (22)
$\leq 0.0625|g_{ji}a_{ji}h^{s}_{i}|+0.25|k_{ji}a_{ji}e_{li}|+1=\max_{y}\left|\partialderivative{\dot{h}^{s}_{i}}{y_{j}}\right|$ (23)
@@
$\left|\partialderivative{\dot{{h^{r}_{i}}}}{y_{j}}\right|$ $=\left|\epsilon(w_{i})\cdot\partialderivative{\dot{h}^{s}_{i}}{y_{j}}+\partialderivative{\epsilon(w_{i})}{y_{j}}\cdot\dot{h}^{s}_{i}\right|,$ (26)
@@
$\leq\epsilon(w_{i})\cdot\left|\partialderivative{\dot{h}^{s}_{i}}{y_{j}}\right|+\epsilon^{\prime}(w_{i})\cdot|o_{ji}|\cdot|\dot{h}^{s}_{i}|$ (28)
@@
$\max_{y}\left|\partialderivative{\dot{h}^{r}_{i}}{y_{j}}\right|$ $\leq\max_{y}\left|\partialderivative{\dot{h}^{s}_{i}}{y_{j}}\right|$ (36)
@@
$\Rightarrow\lambda^{r}$ $<\lambda^{s}$ (39)

### v4

@@ 4.1 LRCUs on RNNs Benchmarks
To achieve a somewhat similar number of parameters, we use 64 cells for LRCUs, and 100 cells for the others as internal representation. The total number of parameters in a model also depends on the number of inputs and outputs of the tasks considered. Finally, we also conduct experiments in a relatively complex and high-dimensional image-based regression task, for lane keeping in autonomous vehicles. All experimental details, including the hyperparameters and the number of trainable parameters, are given in Appendix 0.A.3. In Table 2 we provide the accuracy of the various models compared in the classification tasks.
Table 2: Accuracy in the classification datasets in percentage.
Model Localization IMDB psMNIST
LSTM $82.90\pm 0.31$ $86.56\pm 0.49$ $91.20\pm 0.10$
GRU $82.76\pm 0.41$ $86.32\pm 0.51$ $90.20\pm 0.44$
MGU $83.35\pm 0.30$ $85.18\pm 0.85$ $87.78\pm 0.82$
LRCU-S $\mathbf{84.21\pm 0.26}$ $85.73\pm 0.41$ $\mathbf{91.74\pm 0.41}$
LRCU-A $83.90\pm 0.34$ $\mathbf{87.00\pm 0.53}$ $91.31\pm 0.33$
@@
The Localization Data for the Person Activity dataset given in [34], captures the recordings of five individuals, engaging in various activities. Each wore four sensors at the left and the right ankle, at the chest, and at the belt, while repeating the same activity five times. The associated task, is to classify their activity based on the irregularly sampled time-series. This task is a classification problem, adapted from [22].
In Table 2, we present the experimental results for the accuracy of the classification for LSTMs, GRUs, MGUs, and LRCUs. The experiments for the LRCUs were done with both the asymmetric (LRCU-A) and the symmetric (LRCU-S) elastance, respectively. As one can see, both LRCUs considerably outperformed the traditional gated RNNs. In particular, the LRCU-S performed best, achieving an accuracy of 84.99%, slightly better than the LRCU-A.
@@
The IMDB Movie-Review Sentiment Classification dataset, also known as the Large Movie Review Dataset [27], is designed for binary sentiment classification. This data set includes 25,000 movie reviews, for both training and testing. Each review was labeled with either a positive or a negative sentiment.
Table 2 presents our experimental results for this task, too. While the LRCU-S achieved a performance which was comparable to the one of the traditional RNNs, the LRCU-A had the best performance, by achieving an accuracy of 87%.
@@
The Permuted Sequential MNIST dataset is a variant of the MNIST Digits Classification data set, designed to evaluate recurrent neural networks, adapted from [21]. In this task, the 784 pixels (originally images of size $28{\times}28$) of digits are presented sequentially to the network. The main challenge lies in predicting the digit category, only after all pixels are observed. This task tests the network’s ability to handle long-range dependencies. To make the task even more complex, a fixed random permutation of the pixels is first applied.
Our experimental results are shown in Table 2. As before, both LRCUs surpass the accuracy of the other models. In particular, the LRCU-S achieves the best results.
Figure 6: Lane-Keeping task. The first two figures show summer and winter conditions, respectively. The red rectangle indicates the input and the blue line the predicted steering angle of the network.
Figure 7: Network architecture for the Lane-Keeping task. The CNN head extracts the input features of a video stream. They are passed to the recurrent policy, responsible for steering.
Lane-Keeping Task.
The dataset for the Lane-Keeping Task task is obtained from human-driving recordings captured under various weather and illumination conditions [23], as shown in Figure 6. The network architecture illustrated in Figure 7 contains a CNN-head for extracting the main features from the camera input. They are fed into the gated recurrent models, for the sequential-regression prediction. This setup is adapted from [12].
In Table 3, we report the experimental losses of the models considered on the Lane-Keeping task. As it was the case before, both LRCU-A and LRCU-S obtain comparable validation losses, with LRCU-A achieving the best weighted-validation loss (loss weighted by road curvature). This is arguably the more important loss.
We test the models in the closed-loop setting to evaluate their performance of robust lane keeping. By injecting extra Gaussian noise with variance 0.1 into the input images, we find that LRCUs can handle it without any crashes, and show better performance than other models in even higher Gaussian noise with variance 0.2 too, as reported in Table 4.
Table 3: Mean squared loss results for the Lane-Keeping task, averaged over 3 seeds.
@@
[…]
Table 4: Crash likelihood in closed-loop simulation with additional Gaussian noise in the Lane-Keeping task, averaged over 3 seeds. Values closer to zero indicate reliable, crash-free behavior.
Model Noise $\sigma^{2}=0.1$ Noise $\sigma^{2}=0.2$
LSTM $\mathbf{0.000\pm 0.000}$ $0.483\pm 0.254$
GRU $0.133\pm 0.189$ $0.767\pm 0.047$
MGU $0.267\pm 0.221$ $0.817\pm 0.037$
LRCU-A $\mathbf{0.000\pm 0.000}$ $\mathbf{0.200\pm 0.231}$
LRCU-S $\mathbf{0.000\pm 0.000}$ $0.300\pm 0.216$
4.2 LRCs on Neural ODEs Benchmarks
LRCs are competitive in accuracy and convergence speed on popular Neural-ODE benchmarks. As examples, we consider a Periodic Sinusoidal, a Spiral towards the origin [7], the non-linear Duffing oscillator [28, 31, 10] and three versions of the Lotka-Volterra (LV) predator-prey model [2, 19] well-known from evolutionary biology and ecology.
In our experiments we use an LRC with 16 hidden states for internal representation. LRCs are highly suitable for such tasks as they are ODEs. We also use two extra layers for mapping. The former maps the state of these systems to the dimensionality of the hidden state, allowing us to learn more expressive 16-dimensional dynamics. The latter remaps these dynamics back to the original system state dimensionality.
We compare the performance of the LRCs with a 3-layer Neural-ODE, containing 32, 32, and 2 neurons, respectively. These three layers match the number of trainable parameters, with the ones of the LRCs. Neural-ODEs are solved by Runge-Kutta 4(5) of Dormand-Prince, and LRCs by explicit Euler of order 1. The results of our experiments averaged over 3 seeds, are shown in Table 5.
Table 5: Test loss of various ODE tasks, with LRCs showing good performance. Results are averaged over 3 seeds.
@@ 4.2 LRCs on Neural ODEs Benchmarks
[…]
@@ 4.3 Interpretability of LRCUs on the Autonomous-Driving Lane-Keeping Task
To show that LRCUs outperform LSTMs, GRUs, and MGUs in terms of interpretability, we conduct Lane-Keeping Task experiments, and compute the following interpretability metrics: the network attention and its robustness by the structural similarity index, the absolute correlation of neural activity with road trajectory, and the activity of neurons.
Understanding neural-network attention during decision-making is a crucial element in increasing trust in the network. By using the VisualBackprop [4] method, we visualize which pixels in the input image have the most impact at the given timestep. Even though the same structure of CNN heads is used in the training pipeline, the different recurrent parts of the decision-making policy, have a big influence on the learned features in the convolutional part through backpropagation. The attention in summer and winter season is given in the Appendix.
By injecting a zero mean Gaussian noise with a variance of 0.1 and 0.2 into the tests, respectively, we can measure how robust their attention is. Quantitatively, this is measured by the Structural Similarity Index, displayed in Figure 8. The LRCU models maintain the most similar focus of their attention in the presence of noise, indicated by the boxplots closer to 1. Lighter color refers to additional Gaussian noise of 0-mean and $\sigma^{2}=0.1$ variance, and darker to $\sigma^{2}=0.2$ variance.
As another interpretability metric, we assess how the neural activity within a Lane-Keeping RNN-policy changes during deployment in the closed-loop simulation. Specifically, we are interested in identifying neurons exhibiting an activity matching the geometry of the trajectory. We conduct tests for a 1 km long drive, in both summer and winter. In Figure 9, we illustrate the activity of States $h_{1}$-$h_{2}$, corresponding to Cells 1-2 in the RNN policy, for summer, respectively. In the Appendix, we illustrate the activity of Cells 1-2 for winter. Our results reveal that the LSTM, GRU, MGU cells have scattered, hard-to-interpret activity, respectively. In contrast, LRCU-S cells demonstrate a very gentle varying activity, that aligns very well with the road’s trajectory. This is most likely due to their double-liquid resistance and capacitance time constant.
Table 6: Interpretability. Absolute correlation between the neural activity and the trajectory of the road. Values closer to 1 indicate stronger correlation. Averaged over 3 runs.
Summer Winter
LSTM $0.400\pm 0.299$ $0.450\pm 0.262$
GRU $0.363\pm 0.256$ $0.389\pm 0.248$
MGU $0.317\pm 0.227$ $0.308\pm 0.232$
LRCU-A $0.500\pm 0.314$ $0.533\pm 0.292$
LRCU-S $\mathbf{0.645\pm 0.321}$ $\mathbf{0.766\pm 0.243}$
Refer to captionRefer to caption
Figure 8: Robustness of the attention, measured by the Structural Similarity Index (SSIM) [35] of the models, in summer (left) and in winter (right).
@@
Refer to caption Refer to caption
@@
Refer to caption Refer to caption
Figure 9: Interpretability. The neural activity of two command cells in the learned policy for the Lane-Keeping Task, projected over time on the 1km road driven in summer.
@@ 5 Discussion, Scope and Conclusion
We introduced Liquid-Resistance Liquid-Capacitance networks (LRCs), a Neural ODE model which is more accurate, more efficient, and biologically more plausible, compared to neuroscience’s Electric Equivalent Circuits, and (saturated) Liquid Time-Constant networks (STCs and LTCs).
Due to their gentle varying behavior LRCs are easy to integrate and applicable to a wide range of scenarios, including continuous processes and datasets sampled regularly or irregularly. LRCs thus represent a promising advancement in the field of bio-inspired deep learning. LRCs integrated with explicit Euler of order 1, and a time-step of 1 (LRCUs), have a very similar structure to gated RNNs. They not only demonstrate comparable performance to LSTMs, GRUs, and MGUs on several benchmarks, but also exhibit significantly higher interpretability, demonstrated in the autonomous Lane-Keeping task.
Despite the limitation of the benchmarks used, which we acknowledge could be expanded in subsequent work, our focus on analyzing the properties of our advanced bio-inspired model and its performance across diverse scenarios demonstrates its potential and lays a solid foundation for future exploration. Our results suggest that incorporating additional concepts from neuroscience can advance machine learning models, offering opportunities for future research to explore and further enhance model performance.
Appendix 0.A Appendix
This appendix provides the proofs of the two theorems stated for LRCs, discusses the RNN form of LRCUs and their connection to GRUs and other gated recurrent units, and gives additional details for the experimental evaluation, demonstrating the accuracy, efficiency, and interpretability of LRCs.
0.A.1 Proofs of the Theorems
In this section we provide a detailed proof of the Theorems stated in this paper. The LRC Lipschitz Theorem first derives the local Lipschitz constant of LRCs and then shows that it is smaller than the one of STCs. The LRC Generalization Theorem, first computes the error of LRCs on the validation time series, and then shows that it is less than the one of the STCs, considering the validation set being drawn from in-distribution.
0.A.1.1 Proof of the Lipschitz constant Theorem
Theorem 0.A.1 (LRC Lipschitz).
@@ 0.A.1.1 Proof of the Lipschitz constant Theorem
We now want to prove that there exists a choice of parameters for $\epsilon(w)$, such that the Lipschitz constant of the output $o^{r}$ of the LRC with respect to its hidden states and input $y=[h,x]$, is less than the one of $o^{s}$ in the STC. So the goal is to state conditions on the parameters of $\epsilon(w)$ such that:
$\max_{y}\|\nabla_{y}o^{s}\|_{2}\leq\lambda^{s}\ \wedge\max_{y}\|\nabla_{y}o^{r}\|_{2}\leq\lambda^{r}\quad\Rightarrow\quad\lambda^{r}<\lambda^{s}.$ (6)
@@
$o=Qh,\quad$ $\quad o_{k}=\sum_{i=1}^{m}q_{ki}h_{i},$ (7)
$\partialderivative{o_{k}}{y_{j}}$ $=\sum_{i=1}^{m}q_{ki}\partialderivative{h_{i}}{y_{j}}.$ (8)
@@
$\left|\partialderivative{h_{t}}{y_{t,j}}\right|$ $=\left|\partialderivative{h_{t-\Delta t}}{y_{t,j}}+\partialderivative{y_{t,j}}\int_{t-\Delta t}^{t}\dot{h}dt\right|$ (9)
$=\left|\underbrace{\partialderivative{h_{t-\Delta t}}{y_{t,j}}}_{\leq 1}+\int_{t-\Delta t}^{t}\partialderivative{\dot{h}}{y_{t,j}}dt\right|$ (10)
$\leq 1+\left|\int_{t-\Delta t}^{t}\max_{y}\partialderivative{\dot{h}}{y_{t,j}}dt\right|.$ (11)
@@
$\sigma^{\prime}(x)\leq 0.25,$ $\quad\tau^{\prime}(x)\leq 1.$ (15)
The derivative of the sigmoid function $\sigma^{\prime}(x)=\sigma(x)(1-\sigma(x))$ has a known upper bound of 0.25, which occurs when x = 0. The derivative of the hyperbolic tangent function, $\tau^{\prime}(x)=\mathrm{sech}^{2}(x)$, has a known upper bound of 1, which occurs when x = 0. Using these bounds, for the symmetric version of elastance in Equation (5) it holds that:
$\epsilon^{\prime}(w_{i})=\sigma^{\prime}(w_{i}+k_{i})-\sigma^{\prime}(w_{i}-k_{i})$ (16)
$\leq\max\sigma^{\prime}(w_{i}+k_{i})-\min\sigma^{\prime}(w_{i}-k_{i})\leq 0.25.$ (17)
@@
$\frac{\epsilon^{\prime}(w_{i})}{1-\epsilon(w_{i})}\leq\epsilon(w_{i}).$ (18)
For asymmetric $\epsilon$ this is straightforward as the derivative of a sigmoid is $\sigma^{\prime}(x)=\sigma(x)\cdot(1-\sigma(x))$. For symmetric $\epsilon$ we need to reformulate the equations using that $\sigma(w_{i}+k_{i})\geq\sigma(w_{i}-k_{i})$ for $k_{i}\geq 0$:
$\frac{\epsilon^{\prime}(w_{i})}{1-\epsilon(w_{i})}=$ $\frac{\sigma(w_{i}+k_{i})(1-\sigma(w_{i}+k_{i}))}{1-\sigma(w_{i}+k_{i})+\sigma(w_{i}-k_{i})}$ (19)
$-\frac{\sigma(w_{i}-k_{i})(1-\sigma(w_{i}-k_{i}))}{1-\sigma(w_{i}+k_{i})+\sigma(w_{i}-k_{i})}$ (20)
$\leq\frac{\sigma(w_{i}+k_{i})(1-\sigma(w_{i}+k_{i}))}{1-\sigma(w_{i}+k_{i})}$ (21)
$-\frac{\sigma(w_{i}-k_{i})(1-\sigma(w_{i}+k_{i}))}{1-\sigma(w_{i}+k_{i})}$ (22)
$=\epsilon(w_{i}).$ (23)
@@
$\leq|\sigma^{\prime}(f_{i})g_{ji}\sigma^{\prime}(a_{ji}y_{j}+b_{ji})a_{ji}h^{s}_{i}|$ (25)
$+|\tau^{\prime}(u_{i})k_{ji}\sigma^{\prime}(a_{ji}y_{j}+b_{ji})a_{ji}e_{li}|+1$ (26)
$\leq 0.0625|g_{ji}a_{ji}h^{s}_{i}|+0.25|k_{ji}a_{ji}e_{li}|+1$ (27)
$=\max_{y}\left|\partialderivative{\dot{h}^{s}_{i}}{y_{j}}\right|.$ (28)
@@
$\left|\partialderivative{\dot{{h^{r}_{i}}}}{y_{j}}\right|$ $=\left|\epsilon(w_{i})\cdot\partialderivative{\dot{h}^{s}_{i}}{y_{j}}+\partialderivative{\epsilon(w_{i})}{y_{j}}\cdot\dot{h}^{s}_{i}\right|.$ (31)
@@
$\leq\epsilon(w_{i})\cdot\left|\partialderivative{\dot{h}^{s}_{i}}{y_{j}}\right|+\epsilon^{\prime}(w_{i})\cdot|o_{ji}|\cdot|\dot{h}^{s}_{i}|.$ (33)
@@
$\max_{y}\left|\partialderivative{\dot{h}^{r}_{i}}{y_{j}}\right|\leq$ $\max_{y}\left|\partialderivative{\dot{h}^{s}_{i}}{y_{j}}\right|.$ (43)
@@
$\Rightarrow\lambda^{r}$ $<\lambda^{s}.$ (46)